# Exploring the mysterious high-ionization source powering [NeĀ V] in high- analog SBS0335-052Ā E with JWST/MIRI

### v1

@@ 5 Discussion
S3-Pa | (?) | ā | 7.38 | Bright H lines (T09) | |
S7 | (?) | ā | 7.10 | Bright H lines (T09); [NeĀ V]/[NeĀ II]Ā and [OĀ IV]/[NeĀ III]Ā peak | |
S8 | ā | ā | 7.21 | ||
@@ 5.1 The bright point source, SSC1
As introduced in Sec.Ā 3, the point source is the source SSC1 (Fig.Ā 1). SSC1 appears as a point source in the MIRI/MRS data possibly because of its dusty nature, which makes it particularly bright in the MIR. As described in Sec.Ā 3.2, when analyzing the MIR continuum, we find a very high attenuation for the point source spectrum (). Indeed, as shown in Fig.Ā 2, this spectrum is characterized by the silicate feature at 9.7Ā m, which is much fainter in the other regions of the galaxy. This is consistent with the fact that Hunt etĀ al. (2014) revealed dust continuum emission uniquely coming from the region around SSC1 using ALMA data (see their Fig.Ā 1). This region is also uniquely characterized by emission, tracing warm molecular gas, as first revealed by NICMOS data (Reines etĀ al. 2008; Thompson etĀ al. 2009). We confirm this with our data, observing several H2 rotational transitions around SSC1 (from S(1) to S(7)), which will be modeled and discussed in detail in a forthcoming paper (del Valle-Espinosa et al., in prep.). However, the emission may come from a more extended region, given that the H2 spatial distributions do not show the PSF pattern (see also Thompson etĀ al. 2009; their Fig.Ā 3). The point source region and close surroundings also show high HĀ and Pa equivalent widths (Reines etĀ al. 2008) and purely thermal radio free-free emission, which has been interpreted as due to active star-formation in a dense environment (i.e., Ā O7.5Ā V stars, Ā cm-3, 1.3Ā Mā/yr-1; Johnson etĀ al. 2009).
@@ 5.2 Origin of the highly ionized extended emission
The difficulty in accounting for SBSĀ 0335-052Ā Eās extended high ionization emission was first described by Herenz etĀ al. (2017), who showed a nebular HeĀ IIĀ 4686 map from MUSE that brightened towards the north-west at the rim of a starburst driven super-shell (see also Kehrig etĀ al. 2018), the HeĀ IIĀ shell that is out of our MIRI/MRS FOV. In the following sections, we discuss in detail the different scenarios which can explain the sources of the extended MIR emission, drawing insights from previous studies. In particular, it is important to consider that [OĀ IV]Ā has a similar I.P. to HeĀ IIĀ (Ā ryd), while [NeĀ V]Ā requires the presence of even harder radiation (Ā ryd), corresponding to the extreme ultraviolet and soft X-ray range.
@@ 5.2.1 Pure Bursts of Star Formation
As mentioned previously in Sec.Ā 5.1, Kehrig etĀ al. (2018) was only able to derive SBSĀ 0335-052Ā Eās HeĀ II-ionizing budget via either single, rotating metal-free massive stars (Ā MĀ Ā Mā stars, see their Sec.Ā 5.3.1) or a binary population with 0.05%Ā Zā metallicity and a top-heavy IMF (modeled with BPASS; see their Sec.Ā 5.3.2). However, a 0-0.05%Ā Zā metallicity is unphysically low for SBSĀ 0335-052Ā E (see also Wofford etĀ al. 2021), making this scenario unlikely. Kehrig etĀ al. (2018) discarded also Wolf-Rayet (WR) stars - massive stars (Ā Mā) with strong winds (e.g., Schaerer & Vacca 1998) - as the main source of ionization. Despite WR features being observed in this galaxy (in S3 and in the south-west region of the galaxy, below S8; Papaderos etĀ al. 2006; Izotov etĀ al. 2006; Kehrig etĀ al. 2018), the large number needed to justify the HeĀ II-ionization budget () would overcome by a factor of the SSCs total mass, depending on the IMF choice (Kehrig etĀ al. 2018). The same argument allows us to discard WRs as the main source of [OĀ IV]Ā and thus [NeĀ V]Ā emission.
Wofford etĀ al. (2021; see their SecĀ 3.1) proposed the presence of very massive stars (VMSs) - stars with masses above 100Ā Mā (Vink 2012) - from the blue-shifted [OĀ V]Ā 1371 detection in HST COS data, covering a 2.5ā circular region centered on the peak of the UV emission (see Fig.Ā 1). Thus, one might suspect that VMSs might represent the definite answer, hardening the radiation field and enabling highly ionized extended emission. However, VMSs, while hot and luminous, are known to have dense winds which become optically thick at the short wavelengths needed to power HeĀ IIĀ and [NeĀ V]Ā ionizing photons, as demonstrated by Sander (2022; see also Hawcroft et al. 2025, submitted). Although, a mechanism for the enhanced escape of the hard ionizing radiation from these winds has recently been proposed by Roy etĀ al. (2025), whereby the winds of massive stars are inhomogeneous due to intrinsic instabilities of the outflow. Roy etĀ al. (2025) propose that the higher the degree of inhomogeneity in the outflow, the higher the emergence of extreme UV fluxes. While this possibility has not yet been comprehensively evaluated, it may motivate revisions of stellar population models in the future.
@@ 5.2.3 Shocks
Shocks models are generally considered an important source of ionization in MIR diagnostic diagrams and can be due to several phenomena, including cloud-cloud collisions, the expansion of HĀ IIĀ regions, outflows and/or supernovae (e.g., Groves etĀ al. 2004). Given the disturbed kinematic of this galaxy (Fig.Ā 7; see also Izotov etĀ al. 2006; Herenz etĀ al. 2023), the presence of clear shells in the HĀ and HeĀ IIĀ morphology (Herenz etĀ al. 2017; Kehrig etĀ al. 2018) as well as diffuse non-thermal radio emission interpreted as an ensemble of compact supernova remnants expanding in a dense ISM (Hunt etĀ al. 2004), it is plausible that shocks are present and can play an important role. Unfortunately, the shell structure clearly visible in the HeĀ IIĀ morphology is not covered in our MIRI/MRS observations. This region is slightly north-west to the area where we see Ā km/s red-shifted velocities and intrinsic velocity dispersion enhancement in [SĀ IV]Ā and [NeĀ III]Ā (Fig.Ā 7).
Having said that, as shown in Figs.Ā 8, 13 and 14, low-metallicity F24 (and also AM19, not shown) shock models do not overlap with our data. The grids could match the range of observed [OĀ IV]/[NeĀ III], but cannot explain the corresponding [NeĀ V]/[NeĀ II], [NeĀ III]/[NeĀ II]Ā and [SĀ IV]/[NeĀ II]Ā line ratios. Another hint towards shocks not being the primary source of ionization of the extended [OĀ IV]Ā and [NeĀ V]Ā emission is the complete lack of co-spatial [FeĀ II]Ā emission (see Sec.Ā 5.1). Finally, Kehrig etĀ al. (2018) also excluded shocks as the cause of the HeĀ IIĀ emission, given the lack of high ([OĀ I]Ā 6300) and lack of low [SĀ II]/HĀ line ratios, generally considered to be good shock indicators (e.g., Kewley etĀ al. 2019; Mingozzi etĀ al. 2024).
@@ 5.2.4 The possibility of an IMBH
Recent works have suggested the presence of a Ā Mā IMBH in SBSĀ 0335-052Ā E, due to the detection of [NeĀ V]Ā 3426 (Hatano etĀ al. 2024), as well as possible variability in NEOWISE data (Ā m) over 12 years (Hatano etĀ al. 2023). Interestingly, the [NeĀ V]Ā 3426 was first observed by Thuan & Izotov 2005 and does not show any sign of variability (Hatano etĀ al. 2024). Hatano etĀ al. (2023) proposed also the presence of broad HĀ emission possibly due to the BH broad line region, from which they estimate a Ā Mā upper limit. However, this value is much larger than the mass of the different components of this galaxy (see Tab.Ā 2), raising doubts that this emission is tracing the BH broad line region (see also Sec.Ā 5.3), unless the system is not virialized (e.g., Bertemes etĀ al. 2025). As such, the existence of [NeĀ V]Ā 3426 and the observed NIR variability seem the two most feasible clues from the literature to suggest the presence of an IMBH in SBSĀ 0335-052Ā E.
@@
In consideration of these suggestions of an IMBH, looking at Fig.Ā 8 and Fig.Ā 13, the models that uniquely cover the observed [NeĀ V]/[NeĀ II], [NeĀ III]/[NeĀ II]Ā and [SĀ IV]/[NeĀ II]Ā line ratios that we measure in SBSĀ 0335-052Ā E are the Ā Mā IMBH models from R22 with an AGN fraction 4-8-16% and log() between and (the closed geometry model can better reach the highest [NeĀ III]/[NeĀ II]). According to the R22 models, we can exclude the possibility of a BH with Ā Mā because this BH mass would imply too low log([NeĀ III]/[NeĀ II]) () and log([NeĀ V]/[NeĀ II]) () with respect to the observed data, regardless of the other parameter selection (i.e., AGN fraction, geometry, gas density, stellar age). We can also exclude R22 Ā Mā grids, which still cannot reach the log([NeĀ V]/[NeĀ II]) values that we observe (). R22 Ā Mā can overlap with the observed [NeĀ V]/[NeĀ II]Ā vs [NeĀ III]/[NeĀ II]Ā at 8% AGN fraction, but at lower/higher fractions, they predict larger/lower [NeĀ III]/[NeĀ II]Ā than what we observe.
@@
A great advantage of the MIRI/MRS data is that we can spatially resolve line ratios and trace the ionization hardness (e.g., [NeĀ V]/[NeĀ II], [OĀ IV]/[NeĀ III]), finding their peak around/north SSCs 4,5 and S7 and with decreasing values towards south-east (see Fig.Ā 6). As already discussed in Sec.Ā 4.3, this peak region is characterized by the most disturbed kinematics with the presence of gas flows (Fig.Ā 7; see also Izotov etĀ al. 2006; Herenz etĀ al. 2023). In particular, the [SĀ IV], [NeĀ III]Ā and [NeĀ V]Ā intrinsic velocity dispersion maps revealed an interesting āV shapeā in the north-west region, starting at the position of SSCs 4-5, high-resolution optical IFU data revealed doubled-peak HĀ and HeĀ IIĀ profiles, suggesting the presence of an expanding shell (Izotov etĀ al. 2006). In this scenario, SSC5 may correspond to the location of the IMBH. Another potentially relevant element is that SSC5 is also found to have a very compact (Ā pc) NIR emitting region not resolved in HST NICMOS (high resolution, 0.075Ā ā/px) data (Thompson etĀ al. 2009).
@@
However, we do stress that even the R22 models do not perfectly overlap with our data, since they are not able to reproduce the combination of highest [OĀ IV]/[NeĀ III]Ā and [NeĀ III]/[NeĀ II]Ā values (see Fig.Ā 14). Many factors could be at play here with regards to the assumptions and parameters held within the models - each of which are difficult to constrain with observations alone due to the degeneracies between them. For instance, R22 predictions can vary as a function of the considered SEDs (R22 considered two extreme values, qso or disk-plaw) and assumptions on accretion rate and spin (incorporated as fixed parameters in the qso SED), as well as the choice of geometry (open vs closed). Also, we stress that [OĀ IV]/[NeĀ III]Ā can be affected by the wavelength-dependent MIRI/MRS PSF variation correction, as well as residual fringing (see Sec.Ā 4.4).
@@ 5.3 Implications for high- galaxies and UV-optical diagnostics
As previously mentioned, one of the most popular criteria to identify AGN activity so far has been via broad permitted emission lines (e.g., H). From a visual inspection of VLT/MUSE observations of SBS 0335-052 E MUSE, it appears that a broad H profile exists at both the position of cluster SSC1 (i.e., the MIR point source with no clear [Ne V] detection) and at the peak of He II, [Ne V], and UV emission, with FWHM up to  km/s (see Fig. 15 and App. C). In particular, both the H and [O III] 5007 emission lines require up to 3-4 Gaussian components to be well fitted, with also [O III] 5007 showing a fainter broad (up to FWHM  km/s) component. Interestingly, broad H and the need of components to reproduce the kinematics of the ionized gas has also been noticed in other nearby BCDs (e.g., James et al. 2009; del Valle-Espinosa et al. 2023; Komarova et al. 2021) and other high- analogs (e.g., Green Peas; Amor“in et al. 2024), with interpretations ranging from turbulent mixing layers, high-density gas, and radiatively driven winds. Despite the origin of this broadening still not being completely understood in these systems, there could be a connection with (some of) the broad-line galaxies revealed at high-.
@@
Exploiting the use of high- analogs, Mingozzi etĀ al. (2024) used optical and UV spectra from the Cosmic Origins Spectrograph (COS) Legacy Archive Spectroscopic SurveY (CLASSY; Berg etĀ al. 2022; James etĀ al. 2022) to provide a toolkit to distinguish star formation and AGN activity at high-. In particular, CLASSY is a treasury of 45 nearby () star-forming galaxies, including SBSĀ 0335-052Ā E, generally characterized by lower stellar masses, higher SFRs and more extreme ionization fields than objects, as typically observed in the EoR. Specifically, Mingozzi etĀ al. (2024) proposed the CĀ III]Ā 1907,9/HeĀ IIĀ 1640 versus OĀ III]Ā 1666/HeĀ IIĀ 1640 as the most reliable UV diagnostic diagram (their Fig.Ā 6), being capable of separating the SF, AGN and shocks at low-metallicity. We re-propose this diagnostic diagram in Fig.Ā 9, showing the CLASSY sample compared with R22 models with the same parameters presented in FigureĀ 8 and discussed in SectionĀ 5.2.4. The fact that SBSĀ 0335-052Ā E (light-pink square) lies very close to the line separators was interpreted by Mingozzi etĀ al. (2024) as being due a combination of its low-metallicity environment and hard stellar radiation field. However, the R22 Ā Mā IMBH model with 8%Ā AGN fraction at the lowest metallicity and highest log() would also explain the UV line ratios observed in SBSĀ 0335-052Ā E. Since the UV emission line fluxes are from spectra integrated over the COS aperture (2.5ā; see Fig.Ā 1) covering the [NeĀ V]Ā 14.3 emitting region, the fact that the UV line ratios can be reproduced by the same set of models that explain part of the spatially resolved MIR line ratios indeed strengthens the validity of the R22 models in describing this object. This finding also highlights a previously unknown degeneracy of this diagnostic diagram in distinguishing between pure star-forming objects and objects with an AGN fraction Ā % (when considering R22 models).
@@
The MIR wavelength range is not observable in high- targets, but JWST does provide coverage of [NeĀ V]Ā in the optical wavelength range (e.g., Cleri etĀ al. 2023; Scholtz etĀ al. 2023; Chisholm etĀ al. 2024). As mentioned at the beginning of Sec.Ā 5, optical [NeĀ V]Ā has been revealed in a few nearby BCDs, including SBSĀ 0335-052Ā E. Thus, it is interesting to see how SBSĀ 0335-052Ā E compares to other BCDs and high- targets within which this emission has been observed. In Fig.Ā 10 we show the diagnostic diagram HeĀ IIĀ 4686/HĀ vs [NeĀ V]Ā /[NeĀ III]Ā 3868 proposed by Chisholm etĀ al. (2024) to identify the presence of accreting IMBHs. Indeed, similarly to the MIR line ratios explored in Fig.Ā 8, both [NeĀ V]/[NeĀ III]Ā and HeĀ II/HĀ are relatively insensitive to metallicity and instead trace the ionization parameter and the radiation hardness (Shirazi & Brinchmann 2012; Cleri etĀ al. 2023; Chisholm etĀ al. 2024). Consistently with the MIR (Fig.Ā 8) and UV line ratios (Fig.Ā 9), SBSĀ 0335-052Ā Eās optical line ratios are also consistent with the R22 Ā Mā IMBH models with 8% AGN fraction. Interestingly, SBSĀ 0335-052Ā Eās [NeĀ V]/[NeĀ III]Ā flux ratio (pink square) is below the median values found for the 13 currently known nearby BCD [NeĀ V]Ā emitters (Izotov etĀ al. 2004; 2012; 2021; Thuan & Izotov 2005) and is of the value found for the galaxy studied by Chisholm etĀ al. (2024).

### v2

@@ 5 Discussion
| S3-Pa | Ā (T09) | 2.3Ā (T09) | 7.38 | Bright H lines (T09) | |
| S7 | Ā (T09) | 9.4Ā (T09) | 7.10 | Bright H lines (T09); [NeĀ V]/[NeĀ II]Ā and [OĀ IV]/[NeĀ III]Ā peak | |
| S8 | Ā (T09) | 2.2Ā (T09) | 7.21 | ||
@@ 5.1 The bright point source, SSC1
As introduced in Sec.Ā 3, the point source is the source SSC1 (Fig.Ā 1). SSC1 appears as a point source in the MIRI/MRS data, possibly because of its dusty nature, which makes it particularly bright in the MIR. As described in Sec.Ā 3.2, when analyzing the MIR continuum, we find a very high attenuation for the point source spectrum (). Indeed, as shown in Fig.Ā 2, this spectrum is characterized by the silicate feature at 9.7Ā m, which is much fainter in the other regions of the galaxy. This is consistent with the fact that Hunt etĀ al. (2014) revealed dust continuum emission uniquely coming from the region around SSC1 using ALMA data (see their Fig.Ā 1). This region is also uniquely characterized by emission, tracing warm molecular gas, as first revealed by NICMOS data (Reines etĀ al. 2008; Thompson etĀ al. 2009). We confirm this with our data, observing several H2 rotational transitions around SSC1 (from S(1) to S(7)), which will be modeled and discussed in detail in a forthcoming paper (del Valle-Espinosa et al., in prep.). However, the emission may come from a more extended region, given that the H2 spatial distributions do not show the PSF pattern (see also Thompson etĀ al. 2009; their Fig.Ā 3). The point source region and close surroundings also show high HĀ and Pa equivalent widths (Reines etĀ al. 2008) and purely thermal radio free-free emission, which has been interpreted as due to active star-formation in a dense environment (i.e., Ā O7.5Ā V stars, Ā cm-3, 1.3Ā Mā/yr-1; Johnson etĀ al. 2009).
@@ 5.2 Origin of the highly ionized extended emission
The difficulty in accounting for SBSĀ 0335-052Ā Eās extended high ionization emission was first described by Herenz etĀ al. (2017), who showed a nebular HeĀ IIĀ 4686 map from MUSE that brightened towards the north-west at the rim of a starburst driven super-shell (see also Kehrig etĀ al. 2018), the HeĀ IIĀ shell that is out of our MIRI/MRS FOV. In the following sections, we discuss in detail the different scenarios which can explain the sources of the extended MIR emission, drawing insights from previous studies. In particular, it is important to consider that [OĀ IV]Ā has a similar I.P. to HeĀ II, while [NeĀ V]Ā requires the presence of even harder radiation, corresponding to the extreme ultraviolet and soft X-ray range (see Tab.Ā 1).
@@ 5.2.1 Pure Bursts of Star Formation
As mentioned previously in Sec.Ā 5.1, Kehrig etĀ al. (2018) was only able to derive SBSĀ 0335-052Ā Eās HeĀ II-ionizing budget via either single, rotating metal-free massive stars (see their Sec.Ā 5.3.1) or a binary population with 0.05%Ā Zā metallicity and a top-heavy IMF (modeled with BPASS; see their Sec.Ā 5.3.2). However, a 0-0.05%Ā Zā metallicity is unphysically low for SBSĀ 0335-052Ā E (see also Wofford etĀ al. 2021), making this scenario unlikely. Kehrig etĀ al. (2018) discarded also Wolf-Rayet (WR) stars - massive stars (Ā Mā) with strong winds (e.g., Schaerer & Vacca 1998) - as the main source of ionization. Despite WR features being observed in this galaxy (in S3 and in the south-west region of the galaxy, below S8; Papaderos etĀ al. 2006; Izotov etĀ al. 2006; Kehrig etĀ al. 2018), the large number needed to justify the HeĀ II-ionization budget would overcome by a factor of the SSCs total mass, depending on the IMF choice (Kehrig etĀ al. 2018). The same argument allows us to discard WRs as the main source of [OĀ IV]Ā and thus [NeĀ V]Ā emission.
Wofford etĀ al. (2021; see their SecĀ 3.1) proposed the presence of very massive stars (VMSs) - stars with masses above 100Ā Mā (Vink 2012) - from the blueshifted [OĀ V]Ā 1371 detection in HST COS data, covering a 2.5ā circular region centered on the peak of the UV emission (see Fig.Ā 1). Thus, one might suspect that VMSs might represent the definite answer, hardening the radiation field and enabling highly ionized extended emission. However, VMSs, while hot and luminous, are known to have dense winds which become optically thick at the short wavelengths needed to power HeĀ IIĀ and [NeĀ V]Ā ionizing photons, as demonstrated by Sander (2022; see also Hawcroft et al. 2025, submitted). Although, a mechanism for the enhanced escape of the hard ionizing radiation from these winds has recently been proposed by Roy etĀ al. (2025), whereby the winds of massive stars are inhomogeneous due to intrinsic instabilities of the outflow. Roy etĀ al. (2025) propose that the higher the degree of inhomogeneity in the outflow, the higher the emergence of extreme UV fluxes. While this possibility has not yet been comprehensively evaluated, it may motivate revisions of stellar population models in the future.
@@ 5.2.3 Shocks
Shocks models are generally considered an important source of ionization in MIR diagnostic diagrams and can be due to several phenomena, including cloud-cloud collisions, the expansion of HĀ IIĀ regions, outflows and/or supernovae (e.g., Groves etĀ al. 2004). Given the disturbed kinematic of this galaxy (Fig.Ā 7; see also Izotov etĀ al. 2006; Herenz etĀ al. 2023), the presence of clear shells in the HĀ and HeĀ IIĀ morphology (Herenz etĀ al. 2017; Kehrig etĀ al. 2018) as well as diffuse non-thermal radio emission interpreted as an ensemble of compact supernova remnants expanding in a dense ISM (Hunt etĀ al. 2004), it is plausible that shocks are present and can play an important role. Unfortunately, the shell structure clearly visible in the HeĀ IIĀ morphology is not covered in our MIRI/MRS observations. This region is slightly north-west to the area where we see Ā km/s redshifted velocities and intrinsic velocity dispersion enhancement in [SĀ IV]Ā and [NeĀ III]Ā (Fig.Ā 7).
Having said that, as shown in Figs.Ā 8, 13 and 14, low-metallicity F24 (and also AM19, not shown) shock models do not overlap with our data. The grids could match the range of observed [OĀ IV]/[NeĀ III]Ā (and other high-ionization line-ratios, such as [NeĀ V]/[NeĀ III]Ā and [OĀ IV]/Pf), but cannot explain the corresponding [NeĀ V]/[NeĀ II], [NeĀ III]/[NeĀ II]Ā and [SĀ IV]/[NeĀ II]Ā line ratios. Overall, shock models seem to overpredict low-ionization lines such as [NeĀ II]. Another hint towards shocks not being the primary source of ionization of the extended [OĀ IV]Ā and [NeĀ V]Ā emission is the complete lack of co-spatial [FeĀ II]Ā emission (see Sec.Ā 5.1). Finally, Kehrig etĀ al. (2018) also excluded shocks as the cause of the HeĀ IIĀ emission, given the lack of high ([OĀ I]Ā 6300) and lack of low [SĀ II]/HĀ line ratios, generally considered to be good shock indicators (e.g., Kewley etĀ al. 2019; Mingozzi etĀ al. 2024).
@@ 5.2.4 The possibility of an IMBH
Recent works have suggested the presence of a Ā Mā IMBH in SBSĀ 0335-052Ā E, due to the detection of [NeĀ V]Ā 3426 (Hatano etĀ al. 2024), as well as possible variability in NEOWISE data (Ā m) over 12 years (Hatano etĀ al. 2023). Interestingly, the [NeĀ V]Ā 3426 was first observed by Thuan & Izotov 2005 and does not show any sign of variability (Hatano etĀ al. 2024). Hatano etĀ al. (2023) also proposed the presence of broad HĀ emission possibly due to the BH broad line region, from which they estimate a Ā Mā upper limit. However, this value is much larger than the mass of the different components of this galaxy (see Tab.Ā 2), raising doubts that this emission is tracing the BH broad line region (see also Sec.Ā 5.3), unless the system is not virialized (e.g., Bertemes etĀ al. 2025). As such, the existence of [NeĀ V]Ā 3426 and the observed NIR variability seem the two most feasible clues from the literature to suggest the presence of an IMBH in SBSĀ 0335-052Ā E.
@@
In consideration of these suggestions of an IMBH, looking at Fig.Ā 8 and Fig.Ā 13, the models that uniquely cover the observed [NeĀ V]/[NeĀ II], [NeĀ III]/[NeĀ II]Ā and [SĀ IV]/[NeĀ II]Ā line ratios that we measure in SBSĀ 0335-052Ā E are the Ā Mā IMBH models from R22 with an AGN fraction 4-8% and log() between and (the closed geometry model can better reach the highest [NeĀ III]/[NeĀ II]). According to the R22 models, we can exclude the possibility of a BH with Ā Mā because this BH mass would imply too low log([NeĀ III]/[NeĀ II]) () and log([NeĀ V]/[NeĀ II]) () with respect to the observed data, regardless of the other parameter selection (i.e., AGN fraction, geometry, gas density, stellar age). We can also exclude R22 Ā Mā grids, which still cannot reach the log([NeĀ V]/[NeĀ II]) values that we observe (). R22 Ā Mā can overlap with the observed [NeĀ V]/[NeĀ II]Ā vs [NeĀ III]/[NeĀ II]Ā at 8% AGN fraction, but at lower/higher fractions, they predict larger/lower [NeĀ III]/[NeĀ II]Ā than what we observe.
@@
A great advantage of the MIRI/MRS data is that we can spatially resolve line ratios and trace the ionization hardness (e.g., [NeĀ V]/[NeĀ II], [OĀ IV]/[NeĀ III]), finding their peak around/north SSCs 4,5 and S7 and with decreasing values towards south-east (see Fig.Ā 6). As already discussed in Sec.Ā 4.3, this peak region is characterized by the most disturbed kinematics with the presence of gas flows (Fig.Ā 7; see also Izotov etĀ al. 2006; Herenz etĀ al. 2023). In particular, the [SĀ IV], [NeĀ III]Ā and [NeĀ V]Ā velocity dispersion (corrected for instrumental broadening) maps revealed an interesting āV shapeā in the north-west region, starting at the position of SSCs 4-5, high-resolution optical IFU data revealed doubled-peak HĀ and HeĀ IIĀ profiles, suggesting the presence of an expanding shell (Izotov etĀ al. 2006). In this scenario, SSC5 may correspond to the location of the IMBH. Another potentially relevant element is that SSC5 is also found to have a very compact (Ā pc) NIR emitting region not resolved in HST NICMOS (high resolution, 0.075Ā ā/px) data (Thompson etĀ al. 2009).
@@
However, we do stress that even the R22 models do not perfectly overlap with our data, since they are not able to reproduce the combination of highest [OĀ IV]/[NeĀ III]Ā and [NeĀ III]/[NeĀ II]Ā values (see Fig.Ā 14). Many factors could be at play here with regard to the assumptions and parameters held within the models - each of which are difficult to constrain with emission-line ratio diagrams alone due to the degeneracies between them. For instance, R22 predictions can vary as a function of the considered SEDs (R22 considered two extreme values, qso or disk-plaw) and assumptions on accretion rate and spin (incorporated as fixed parameters in the qso SED), as well as the choice of geometry (open vs closed). Also, we stress that [OĀ IV]/[NeĀ III]Ā can be affected by the wavelength-dependent MIRI/MRS PSF variation correction, as well as residual fringing (see Sec.Ā 4.4).
@@ 5.3 Implications for high- galaxies and UV-optical diagnostics
As previously mentioned, one of the most popular criteria to identify AGN activity so far has been via broad permitted emission lines (e.g., H). From a visual inspection of VLT/MUSE observations of SBS 0335-052 E MUSE, it appears that a broad H profile exists at both the position of cluster SSC1 (i.e., the MIR point source with no clear [Ne V] detection) and at the peak of He II, [Ne V], and UV emission, with FWHM up to  km/s (see Fig. 15 and App. C). This implies that it may have an extended nature, not compatible with a BH broad line region. In particular, both the H and [O III] 5007 emission lines require up to 3-4 Gaussian components to be well fitted, with also [O III] 5007 showing a fainter broad (up to FWHM  km/s) component. Interestingly, broad H and the need of components to reproduce the kinematics of the ionized gas has also been noticed in other nearby BCDs (e.g., James et al. 2009; del Valle-Espinosa et al. 2023; Komarova et al. 2021) and other high- analogs (e.g., Green Peas; Amor“in et al. 2024), with interpretations ranging from turbulent mixing layers, high-density gas, and radiatively driven winds. Despite the origin of this broadening still not being completely understood in these systems, there could be a connection with (some of) the broad-line galaxies revealed at high-.
@@
Exploiting the use of high- analogs, Mingozzi etĀ al. (2024) used narrow lines in optical and UV spectra from the Cosmic Origins Spectrograph (COS) Legacy Archive Spectroscopic SurveY (CLASSY; Berg etĀ al. 2022; James etĀ al. 2022) to provide a toolkit to distinguish star formation and AGN activity at high-. In particular, CLASSY is a treasury of 45 nearby () star-forming galaxies, including SBSĀ 0335-052Ā E, generally characterized by lower stellar masses, higher SFRs and more extreme ionization fields than objects, as typically observed in the EoR. Specifically, Mingozzi etĀ al. (2024) proposed the CĀ III]Ā 1907,9/HeĀ IIĀ 1640 versus OĀ III]Ā 1666/HeĀ IIĀ 1640 as the most reliable UV diagnostic diagram (their Fig.Ā 6; see also Feltre etĀ al. 2016), being capable of separating the SF, AGN and shocks at low-metallicity. We revisit this diagnostic diagram in Fig.Ā 9, showing the CLASSY sample compared with R22 models with the same parameters presented in FigureĀ 8 and discussed in SectionĀ 5.2.4. The fact that SBSĀ 0335-052Ā E (light-green square) lies very close to the line separators was interpreted by Mingozzi etĀ al. (2024) as being due a combination of its low-metallicity environment and hard stellar radiation field. However, the R22 Ā Mā IMBH model with 8%Ā AGN fraction at the lowest metallicity and highest log() would also explain the UV line ratios observed in SBSĀ 0335-052Ā E. Since the UV emission line fluxes are from spectra integrated over the COS aperture (2.5ā; see Fig.Ā 1) covering the [NeĀ V]Ā 14.3 emitting region, the fact that the UV line ratios can be reproduced by the same set of models that explain part of the spatially resolved MIR line ratios indeed strengthens the validity of the R22 models in describing this object. This finding also highlights a previously unknown degeneracy of this diagnostic diagram in distinguishing between pure star-forming objects and objects with an AGN fraction Ā % (when considering R22 models with Ā Mā).
@@
The MIR wavelength range is not observable in high- targets, but JWST does provide coverage of [NeĀ V]Ā in the optical wavelength range (e.g., Cleri etĀ al. 2023; Scholtz etĀ al. 2023; Chisholm etĀ al. 2024). As mentioned at the beginning of Sec.Ā 5, optical [NeĀ V]Ā has been revealed in a few nearby BCDs, including SBSĀ 0335-052Ā E. Thus, it is interesting to see how SBSĀ 0335-052Ā E compares to other BCDs and high- targets within which this emission has been observed. In Fig.Ā 10 we show the diagnostic diagram HeĀ IIĀ 4686/HĀ vs [NeĀ V]Ā /[NeĀ III]Ā 3868 proposed by Chisholm etĀ al. (2024) to identify the presence of accreting IMBHs. Indeed, similarly to the MIR line ratios explored in Fig.Ā 8, both [NeĀ V]/[NeĀ III]Ā and HeĀ II/HĀ are relatively insensitive to metallicity and instead trace the ionization parameter and the radiation hardness (Shirazi & Brinchmann 2012; Cleri etĀ al. 2023; Chisholm etĀ al. 2024). Consistently with the MIR (Fig.Ā 8) and UV line ratios (Fig.Ā 9), SBSĀ 0335-052Ā Eās optical line ratios are also consistent with the R22 Ā Mā IMBH models with 8% AGN fraction. Interestingly, SBSĀ 0335-052Ā Eās [NeĀ V]/[NeĀ III]Ā flux ratio (light-green square) is below the median values found for the 13 currently known nearby BCD [NeĀ V]Ā emitters (Izotov etĀ al. 2004; 2012; 2021; Thuan & Izotov 2005) and is of the value found for the galaxy studied by Chisholm etĀ al. (2024).